# Orca: Browsing at Scale Through User-Driven and AI-Facilitated Orchestration Across Malleable Webpages

### v1

@@ 5.1.1. Pinning webpages for fixed and concentrated views
On the other hand, for any webpage the user wants to view in a fixed manner (e.g., a document they are working on) or highlight, they may pin it to the side (Figure 3.4). Pinned webpages are displayed at the original size, as in a traditional browser, and can be switched among through the Pinned Webpage Bar (Figure 2.a). Unlike the traditional tab bar—where all opened tabs across all tasks are shown, burdening users to organize and relocate them among a large volume of tabs—users may dynamically create the bars with only selected webpages, while others stay open on the canvas.
@@ 5.2. Organize at Scale: Flexibly Organizing Webpages with Different Structures
Manually managing webpages on a freeform canvas and adjusting the viewport can be demanding. Unlike positioning shapes in Figma, where layout is part of the design task itself, this effort can be perceived as an extra burden (Kosch et al., 2023). Meanwhile, a cluttered canvas may make it harder to discover connections, synthesize information, or cross-reference pages. Therefore, Orca offers various ways to help gather and tidy webpages, as well as compose them into different structures and representations.
Orca offers grid and stack organizations of webpages. Users can arrange pages into a grid for a clean layout for quick scanning (Figure 4.a). They can also sort the grid by content or browsing metadata (including time opened or last interacted) to understand relationships and quickly locate webpages. To archive content with minimal canvas space while keeping tabs open, or for pages not currently in use, users can use stacks, where they may flip through the webpages inside (Figure 4.b). Webpages can be dragged in or out of grids and stacks as needed.
Prior work has explored representing multiple webpages as diagrams or nested lists to show relationships among them (Ayers and Stasko, 1995; Chang et al., 2021b). To further support diverse understanding and tasks, Orca explores representing sets of webpages with tables and interactive graphics (Jiang et al., 2023b). Orca displays content from each page in columns of a table. Users can continue adding more columns to extract more information across all pages in the table view. Webpages can also be visualized as interactive diagrams (e.g., bar charts) using Vega-Lite grammar generated on-demand for visualizing a specific aspect of the web content across pages, e.g., when comparing hotel prices (Satyanarayan et al., 2017).
@@ 5.3. Navigate at Scale: Batch Exploration in Depth and Breadth Through Direct Manipulation
The tedious process and the effort required to manage context across multiple threads often discourage users from exploring as much as they would like or fully following and using multiple threads. Therefore, Orca streamlines exploration in both directions of depth and breadth through direct manipulation.
To help users quickly drill down and overview multiple links on a page, Orca gathers links into meaningful groups for Batch Open, such as “Discounted hotels” on a hotel listing page (Figure 6). Users can quickly open these links by dragging the capsule button onto the canvas. Users can also define custom extraction queries, such as “Hyatt with rating above 9”. Orca extracts links to detail pages that match the query (i.e. Hyatt hotels rated above 9.0) and shows the total number found. The user can then drag the capsule or press Return to open the links on the side of the source page.
To expand the exploration in breadth, Orca helps users search for and visit relevant pages based on their current context. After selecting one or more pages, users can click a small handle at the bottom right corner to open the Contextual Expansion menu (Figure 7). Orca analyzes the content of selected webpage(s) and suggests relevant queries. For example, if the user selects a hotel booking page in Paris, Orca may suggest queries like “Find flights to Paris” to complete the trip planning or “Find the same hotel on other platforms” to compare the prices and options. Users may also input customized queries. Users can then drag the handle (like resizing a shape) to control how many results to open, and thus the complexity of the broad exploration, or press Return (or double-click) to let AI decide the number of pages. When a target page cannot be accessed directly, Orca deploys a web automation agent to navigate from a starting point and reach the desired page, e.g., starting from a hotel booking homepage, automatically filling in trip details, and reaching the desired detail page (Section 5.6).
As a trade-off for the zoomed-out overview, each page on the Web Canvas appears smaller to fit within the limited screen space. This means web content is shown at a scale that web designers typically do not intend for users (even with the mobile version).
While our prototype shows that users can still easily read and interact with web content at around one-third scale (on a 14-inch monitor)—fitting up to 9 webpages—it is critical to surface relevant content from each page at a familiar, constant size, regardless of the zoom level of the Web Canvas. More importantly, each page often contains clutter that is unrelated to the user’s current task. Helping users quickly spot relevant content across multiple pages can greatly improve their exploration efficiency.
To support targeted, customized information access, Orca enables extracting content and interactive elements from webpages (Figure 5). Users can select multiple webpages, toggle Page Extraction, and enter a query to pull out relevant information. For example, when comparing hotels, they might ask, “Does it have a pool?” or simply, “Pool”, and Orca uses the LLM to interpret the intent, find answers, and display them in a scannable format. Meanwhile, the source of extracted information is highlighted on the original webpage for the user to inspect when needed. This helps users scan through and filter options in a more customized manner than using the search and filter tools offered by the original websites (Min et al., 2025).
Users can also surface interactive components to perform simple actions across pages directly from the overview. Instead of extracting raw webpage code, the Orca prototype provides buttons, input boxes, and text areas as standardized, reusable UI elements that represent the underlying interactive widgets. This approach enables consistent and reliable UI extraction and generation (Cao et al., 2025). The components remain bi-directionally synced with the original web UI via two mechanisms: content listeners update the extracted UI when the page changes, and web automation agents transmit user actions (such as clicks and inputs) back to the page (Section 5.6).
@@ 5.4. Extract at Scale: View Customization Through Continuous Page Extraction
5.5. Synthesize at Scale: Dynamic Summary and Information Query
Having to maximize the view of webpages one by one and read through each can significantly slow down the exploration and discourage users from opening and gaining information from the large number of webpages that Orca helps create and manage.
On the other hand, AI-driven information retrieval pipelines like Retrieval-Augmented Generation (RAG), or systems like Perplexity, typically provide a one-pass summary over the gathered resources in the end, despite supporting follow-up queries. In contrast, Orca supports an iterative, contextual, and user-driven process of collecting and making sense of sources. As users scout and select webpages, they can dynamically surface intermediate summaries that evolve with their exploration: by toggling the Exposé sidebar, users can view individual summaries for each page or an aggregated summary across the selected set (updated and cached on selection or web content changes). The granular and dynamic summaries facilitate interleaved sensemaking and information gathering.
Users can continue refining their information retrieval by asking follow-up questions about the selected webpages (or pages in the viewport if none are selected). If the answer is not directly available from the current content, Orca deploys automation agents to the relevant pages to further navigate and search for broader context in response to the question (Section 5.6).
Prior approaches to web automation typically involve agents completing entire tasks from start to finish with minimal user involvement, such as generating a full report for a question. However, users often struggle to stay updated with the latest information and progress, and cannot effectively provide intermediate feedback for emerging needs and preferences (Huq et al., 2025). Orca takes a different approach by offering ad hoc and “atomic” automation. Agents act as on-demand components that support specific parts of the workflow, allowing for more flexible, user-driven interaction.
As mentioned in previous sections, automation in Orca supports tasks such as navigating newly opened pages to reach the target page, exploring pages to gather additional information in response to user questions, and carrying out direct operational requests, triggered globally or for one or a few selected webpages. Automation is visualized through virtual cursors—one for each agent—allowing users to track and control their actions in real time (Figure 8).
Web Canvas encodes rich context shared across webpages, both visually for the user to track and as global context for agent utilization. As a result, agents are automatically contextualized without requiring users to enter details manually. For example, if several webpages and a form (to be completed using information from those pages) are placed side by side, selecting all and calling an agent to fill out the form allows it to pull necessary information directly from the neighboring page.
Prior systems based on tabbed browsers typically use a single agent to complete tasks sequentially, often leading to slow progress and poor traceability due to frequent tab switches and disconnected views. In contrast, Orca leverages the Web Canvas as a workspace where multiple processes can run and be visualized in parallel, like dashboards or spreadsheets (where computations are carried out concurrently in space (Dourish, 2017)). For example, to check the distance to the airport for a set of hotels, all detail pages can be automatically opened and organized in a grid structure, where an agent is then deployed to each page to navigate to the right section simultaneously and extract the relevant data to generate the summary.
@@ 5.7. Supporting Diverse Scalable Instructions with Feedforward Prompting
While many at-scale controls across webpages can be performed through direct manipulation on the Web Canvas, Orca also provides a global command bar that enables users to orchestrate malleable webpages using natural language prompts directly (Figure 9).
The bar can be toggled with Control/Command + T or a button, and functions like a traditional address bar (navigating to a URL or searching a query) when Return is pressed directly.
We synthesize the user workflow identified in the problem framing (Section 2) along with the corresponding aspects required to enable browse at scale, to derive four types of supported cross-page requests: create, operate, organize, and query (Table 2).
A common problem with prompting LLM-assisted actions is the black-box nature of interpretation (Subramonyam et al., 2024; Zamfirescu-Pereira et al., 2023; Min and Xia, 2025). Users often do not know what the model can do—or how it will respond—until after the action is performed. With an identified set of cross-page actions, Orca supports feedforward prompting by translating open-ended user input into a list of options clearly describing specific system actions to trigger. Users can then select from these options, which the AI will carry out.
@@ 6. Implementation
The Orca prototype is an Electron app built with React, where each webpage is loaded in its own webview (OpenJS Foundation and Electron contributors, [n. d.]). The Web Canvas interface is implemented using tldraw (tld, [n. d.]). An M4 Max MacBook Pro with 36GB of unified memory handles up to around 80 webpages before freezing. All LLM-based tasks, including web automation and content summarization, are powered by Claude 3.7 Sonnet (20250219). For web automation, we developed custom HTML distillation and agentic pipelines, following common practices in LLM-based web automation (Zhou et al., 2023; He et al., 2024; Müller and Žunič, 2024). As a result, we can reuse the distilled web content for Page Extraction (Appendix B) and support flexible user intervention during the automated process.
Orca is open-sourced [^2] https://github.com/peilingjiang/orca to facilitate further exploration of at-scale web browsing, and to serve as a testbed for advancing parallel agentic web automation pipelines.
@@ 7. Preliminary Evaluation
Evaluating a new browser is challenging due to established user habits with tabbed browsing. A comprehensive understanding of many designs and features requires a longitudinal study with daily use. However, security concerns and the performance of the current prototype make it unsuitable for production and beyond this research’s scope. To assess the usability and effectiveness of the new design concepts and features, we conducted preliminary lab studies. We aim to answer the following research questions:
@@ 7.1. Participants
We recruited 8 participants (4 female, 4 male, aged 21–27) via internal channels at a university. All participants regularly use generative AI tools like ChatGPT. They also have experience with canvas-based graphics editing tools (e.g., Figma, Adobe Illustrator)—four participants (P1, P3, P5, P6) use these tools frequently (more than a few times per week), while the others (P2, P4, P7, P8) use them less often (less than a few times per month).

### v2

@@ 5.1.1. Pinning webpages for fixed and concentrated views
As Leela narrows down hotels, she starts to note them down in Google Docs. She pins this note page to the left sidebar so it stays full size, then continues to rearrange, open, and close hotel pages on the Web Canvas. Whenever she finds a promising option, she glances at the pinned note, updates her shortlist, and hides the sidebar when she wants a distraction-free canvas view.
For any webpage that users want to keep in a fixed, prominent view (for example, a document they are editing or a reference page), they can pin it to the side (Figure 3.4). Pinned webpages are displayed at their original size, as in a traditional browser, and can be switched via the Pinned Webpage Bar (Figure 2.a). Unlike a traditional tab bar—which shows every open tab and forces users to manage them all at once—Orca lets users create dynamic tab bars that hold only selected webpages, while other pages remain open on the Web Canvas.
@@ 5.2. Organize at Scale: Flexibly Organizing Webpages with Different Structures
After twenty minutes of exploring, Leela has dozens of hotel pages scattered across the Web Canvas. She range-selects them and asks Orca to organize them into a grid for a clean overview. After sorting the options based on her preferences, she switches the grid into a stack to set aside before exploring flights.
Manually arranging webpages on a freeform canvas and managing the viewport can be demanding. Unlike positioning shapes in Figma, where layout is part of the design task, such effort can feel like an added burden (Kosch et al., 2023). A cluttered canvas also makes it harder to discover connections, synthesize information, or cross-reference pages. To address this, Orca provides ways to gather, tidy, and recompose webpages into different structures.
Orcaoffers grid and stack organizations. Users can place pages in a grid for a clean, scannable layout (Figure 4.a) and sort the grid by content or browsing metadata (such as time opened or last interacted) to reveal relationships or locate pages quickly. To archive content while keeping tabs available, or for pages not in active use, users can place them in stacks and flip through them as needed (Figure 4.b). Webpages can be dragged in or out of grids and stacks.
Prior work has explored representing multiple webpages as diagrams or nested lists to show their relationships (Ayers and Stasko, 1995; Chang et al., 2021b). Orca explores representing webpages with tables and interactive graphics (Jiang et al., 2023b). Orca displays content from each page in columns of a table. Users can continue adding more columns to extract more information across all pages in the table. Webpages can also be visualized as interactive diagrams (e.g., bar charts) using Vega-Lite grammar generated on demand for visualizing specific aspects of content across pages, e.g., when comparing hotel prices (Satyanarayan et al., 2017).
With the hotel stack set aside, Leela opens a flight search page for her trip dates. Instead of going through the results one by one, she uses Batch Open capsule to filter and open all nonstop flights that match her preferred departure window into a grid. From one flight detail page, she then pulls out the Contextual Expansion handle to explore a few alternative routes from other airlines into the same region of the canvas, dragging the handle to reserve space for just a handful of extra options.
@@ 5.3. Navigate at Scale: Batch Exploration in Depth and Breadth Through Direct Manipulation
This figure contrasts two organizational layouts for a group of webpages. The top image shows a grid layout, with three webpages displayed side-by-side. The bottom image shows a stack layout, where only the first page of the group is visible. Both views feature a shared control bar below the pages that displays the group’s name and page count, along with icons to manage the group. A cursor is shown clicking the button to switch from the grid to the stack layout.
@@
A three-part diagram shows the process of extracting information across multiple webpages. The process begins with a grid of hotel webpages. Next, after the user queries for ‘Bar’, the view changes so each webpage preview is covered by a simple card showing ‘Yes’ or ‘No’. In the final step, another query is input by the user for the ‘Distance to airport’, and the cards in the grid update to show both pieces of information for each hotel. A cursor is shown clicking the button to change the webpage grid to the chart view.
@@
A four-part diagram demonstrates a feature for opening multiple links from a webpage at once. The process starts with a user hovering over the ‘sidecar’ menu on a movie-ranking webpage, minimized to avoid distraction, which enlarges and reveals a list of suggested queries to extract and open links in the page. The user then inputs a custom query of ‘rating above 9.0’, dragging the purple capsule that represents the 7 matching links onto the Web Canvas, which automatically opens all 7 movie pages and arranges them in a grid.
The tedious process and the effort required to manage context across multiple threads often discourage users from exploring or fully utilizing different threads. Therefore, Orca streamlines exploration in both directions of depth and breadth through direct manipulation within Web Canvas.
To help users quickly drill down and preview many links, Orca can Batch Open links, such as “Discounted hotels” on a listing page (Figure 6). Users open these links by dragging the capsule button onto the canvas. They can also define custom extraction queries, such as “Hyatt, rating above 9”. Orca extracts matching detail pages, shows the total number found, and lets users drag the capsule or press Return to open them on the side.
To expand exploration in breadth, Orca suggests relevant pages based on context. After selecting one or more pages, users click the small handle in the bottom-right corner to open the Contextual Expansion menu (Figure 7). Orca analyzes the selection and proposes queries. For example, when the user selects a hotel page, it may suggest “Find flights” or “Find on other websites”. Users can also enter custom queries. They can then drag the handle to control how many results to open or press Return to let AI choose. When a target page cannot be reached directly, Orca deploys a web automation agent to navigate from a starting point (e.g., Google) to the desired page (Section 5.6).
@@
A four-part illustration demonstrates the ‘Contextual Expansion’ feature. The process begins with a user clicking a small handle on a hotel webpage, which opens a menu of suggested queries, like checking prices on other sites. The user then drags the handle to designate space for the new page. In the final step, a new webpage from a competing travel site opens next to the original one. A status bar indicates that an automation agent is navigating the new page to find the requested information.
After shortlisting flights, Leela brings her hotel stack back into a grid to check their distances from the airport where her flight lands. She selects the hotel grid and turns on Page Extraction, adding the prompt “Distance to ORY.” Orca overlays each hotel page with the distance provided by the detail page. She quickly removes hotels that are too far from the airport and keeps a small set of viable options.
Our prototype shows that users can still read and interact with web content at about one-third scale on a 14-inch monitor, fitting up to nine pages at once. However, it is critical to surface key content from each page at a familiar, constant size, regardless of the zoom level on the Web Canvas. Many pages also contain clutter that does not relate to the user’s task. Helping users spot relevant content across pages can greatly improve exploration.
To support focused access, Orca lets users extract content from webpages (Figure 5). Users select pages, toggle Page Extraction, and enter a query to pull out relevant information. For example, when comparing hotels, they might ask “Does it have a pool?” or simply “Pool”. Orca uses the LLM to interpret the intent, find answers on page, and display them in a clear, scannable format. The system also highlights the source on each webpage so users can inspect it when needed. This process allows more flexible filtering than the built-in tools on most sites (Min et al., 2025).
Users can also surface interactive elements to perform simple actions across pages directly from the Web Canvas overview. Instead of extracting raw code, Orca provides standardized UI elements—such as buttons, input boxes, and text areas—that represent the underlying widgets, allowing consistent UI extraction and generation (Cao et al., 2025). The extracted components stay bi-directionally synced with the original pages through two mechanisms: content listeners update the extracted UI when the page changes, and web automation agents send user actions back to the page (Section 5.6).
5.5. Synthesize at Scale: Dynamic Summary and Information Query
Once she has a small set of flights and hotels on the Web Canvas, Leela selects them and opens the Exposé sidebar. Orca shows a brief summary for each option and highlights key trade-offs, such as which combinations fit her budget and commute constraints. She then types a follow-up question asking which pairings best match her schedule. She skims the updated summary while rearranging pages on the canvas.
Maximizing pages one by one and reading each in full can slow exploration. It also discourages users from opening and learning from the many webpages that Orca helps them create and manage.
@@ 5.5. Synthesize at Scale: Dynamic Summary and Information Query
An illustration shows two color-coded automation agents working on webpages in parallel. Each agent is represented by a large virtual cursor on the page and a matching colored control bar below it that states its goal. On the left, a green agent selects a date on a travel site. On the right, the user has clicked the purple agent’s cursor to open a menu of commands, including an option to ‘Take over’. A diagram shows the agent’s virtual cursor snapping to the user’s actual cursor to signify that the user is now in direct control.
AI-driven retrieval (e.g., Retrieval-Augmented Generation) systems like Perplexity usually provide a single summary at the end of a session, despite supporting follow-up queries. In contrast, Orca enables an iterative and user-controlled process of interpreting sources. As users open and select webpages, they can surface intermediate summaries that evolve with their browsing. By toggling the Exposé sidebar, they can view summaries for each page or an aggregated summary across the selected set. Orca updates and caches these summaries whenever the selection or web content changes. These granular, dynamic summaries support interleaved sensemaking and information gathering.
Users can refine their retrieval by asking follow-up questions about the selected pages, or the pages in view when none are selected. If the answer is not available from the current content, Orca deploys automation agents to navigate and search for additional context (Section 5.6).
After settling on her flight, hotel, and a few activities, Leela moves on to finalizing the planning. She selects all relevant pages, invokes automation agents—one per page—and asks them to complete the bookings. Each agent pulls from her user profile and the schedule and details in her notes to make the reservation in parallel. The sidebar summarization updates in real time, showing the status of each booking as it completes.
@@ 5.6. Operate at Scale: On-Demand, Contextualized, and Parallel Web Automation
Three examples show how a global command bar suggests actions based on the user’s context. Each example displays the current webpages on the left and the command bar on the right. First, with a flight search to Paris on screen, the command bar suggests related actions like searching for hotels. Second, when the screen is empty, the command bar suggests actions to find and open new pages based on the user’s query. Third, with two pages open, the command bar suggests an action to find information on one of them; hovering over the suggestion highlights the relevant page.
@@
Prior approaches to web automation typically involve agents completing full tasks end-to-end with no user involvement, such as generating a complete report for a query. However, users often struggle to stay updated on progress and cannot provide timely feedback for emerging needs and preferences (Huq et al., 2025). Orca instead offers ad hoc, “atomic” automation: agents act on demand and support specific parts of a user-driven workflow.
As described earlier, automation in Orca supports tasks such as navigating newly opened pages, exploring pages to gather information in response to user questions, and carrying out direct operational requests, whether globally or on selected webpages. Automation is visualized through virtual cursors—one per agent—so users can track and control actions in real time (Figure 8).
Web Canvasencodes rich shared context across webpages, both visually for the user and as global context for agents. As a result, agents become automatically contextualized without requiring manual input. For example, if several webpages and a form are placed side by side, selecting them and invoking an agent to fill the form allows it to pull the context directly from the neighboring pages.
Systems based on tabbed browsers typically rely on a single agent executing tasks sequentially, resulting in slow progress and poor traceability due to constant tab switching and disconnected views. In contrast, Orca leverages the Web Canvas as a workspace where multiple processes run and are visualized in parallel, similar to dashboards or spreadsheets (where computations occur concurrently in space (Dourish, 2017)). For example, to check the distance to an airport for a set of hotels, all detail pages can be opened and organized into a grid, and agents can be deployed to each page to navigate to the relevant section simultaneously and extract the information.
@@ 5.7. Supporting Diverse Scalable Instructions with Feedforward Prompting
While at-scale controls on the Web Canvas can be performed via direct manipulation, Orca also provides a global command bar for orchestrating malleable webpages using natural language (Figure 9).
The bar, toggled with Command + T or a button, behaves like a traditional address bar for navigation or search when users press Return. Building on the workflow identified in the problem framing (Section 2) and the requirements for browsing at scale, we derive four types of supported cross-page requests: create, operate, organize, and query (Table 2).
A common challenge with prompting LLM-assisted actions is their black-box interpretation (Subramonyam et al., 2024; Zamfirescu-Pereira et al., 2023; Min and Xia, 2025): users often cannot predict what the model can do or how it will respond. With a defined set of cross-page actions, Orca supports feedforward prompting by translating open-ended input into clear, available options for system actions. Users then select an option for AI to execute.
@@ 6. Implementation
The Orca prototype is an Electron app built with React, where each webpage is loaded in its own webview (OpenJS Foundation and Electron contributors, ). The Web Canvas interface is implemented using tldraw (104). All LLM-based tasks, including web automation and content summarization, are powered by Claude 3.7 Sonnet (20250219). For web automation, we developed custom HTML distillation and agentic pipelines, following common practices in LLM-based web automation (Zhou et al., 2023; He et al., 2024; Müller and Žunič, 2024). As a result, we can reuse the distilled web content for Page Extraction (Appendix B) and support flexible user intervention during the automated process.
Displaying many webpages can place heavy loads on the device. An M4 Max MacBook Pro with 36GB of RAM can handle about 80 webpages before freezing, which already exceeds what most users open for a typical task (Chang et al., 2021a; Ma et al., 2023). However, when we apply techniques used in modern browsers, such as background tab pausing [^1] https://support.google.com/chrome/answer/12929150, which stops rendering and scripting for inactive pages while preserving the last rendered content, the system remains functional even after loading 500 tabs in testing. No participants experienced related issues during the evaluation.
Because Electron is Chromium-based, our prototype inherits the default security model of other Chromium browsers (e.g., Chrome) for resisting malicious webpages and managing cached user data. In addition, the model providers we use state that they do not access or train on LLM requests sent through their APIs, which is how Orca communicates with them. Given this setup, we believe security and user privacy are reasonably protected for a research prototype, though alternatives such as locally hosted models could further strengthen privacy guarantees [^2] Such as https://ollama.com.
Orca is open-sourced [^3] https://orca.jiang.pl to facilitate further exploration of at-scale web browsing, and to serve as a testbed for advancing parallel agentic web automation pipelines.
@@ 7. Preliminary Evaluation
Evaluating a new browser is challenging due to established user habits with tabbed browsing. A comprehensive understanding of many designs and features requires a longitudinal study with daily use. However, security concerns make it unsuitable for production and beyond this research’s scope. To sufficiently assess the usability and effectiveness of the new design concepts and features, we conducted preliminary lab studies.
Because conventional tabbed browsing already serves as participants’ de facto baseline from daily use, we did not add a separate control condition and instead relied on their existing experience to judge the usability and perceived benefits of Orca.
We aim to answer the following research questions:
@@ 4. RQ4
  5. RQ5
    How does Orca help users identify and recover from AI and system failures?
@@ 7.1. Participants
We recruited 8 participants (4 female, 4 male, aged 21–27) via internal channels at a university. All participants regularly use generative AI tools like ChatGPT. They have varying experience with canvas-based graphics editing tools (e.g., Figma, Adobe Illustrator)—four participants (P1, P3, P5–6) use these tools often (more than a few times per week), P7–8 use them less than a few times per month, and P2 and P4 use them once every few months or less.